OpenStudy (anonymous):

Solve for n. n/9 < n/2 - 1/3

OpenStudy (anonymous):

let us first try to simplify the right hand side \[{n\over2}-{1\over 3}\]

OpenStudy (anonymous):

i would multiply by \(18\) to get rid of the annoying fractions

OpenStudy (anonymous):

this simplifies to \[{n\times3\over2\times3}-{1\times2\over3\times2}=\]

OpenStudy (anonymous):

I agree with satellitey

OpenStudy (anonymous):

Multiply both sides by 18, re-arrange to bring the n's on one side and then divide both sides so that n is left on one side and a number on the other.

OpenStudy (anonymous):

both ways lead to rome

OpenStudy (anonymous):

i hate fractions ! \[\frac{n}{9}<\frac{n}{2}-\frac{1}{3}\] \[18\left (\frac{n}{9}\right)<18\left (\frac{n}{2}-\frac{1}{3}\right)\]\[2n<9n-6\]

OpenStudy (anonymous):

subtract \(2n\) from both sides to get \[0<7n-6\] add \(6\)\[6<7n\]divide by \(7\)

OpenStudy (anonymous):

6/7

OpenStudy (anonymous):

your answer should be an inequality, not a number

OpenStudy (anonymous):

\[\frac{6}{7}<n\] or \[n>\frac{6}{7}\]

OpenStudy (anonymous):

oh thank you

OpenStudy (anonymous):

yw